From GlueXWiki
Jump to: navigation, search

May 6, 2010 FDC meeting

Tentative Agenda

  1. Production
  2. Electronics
    • update on oscillation problem (Fernando)
    • possible grounding improvement
    • PCB (Kim, Fernando)
    • fADC125 test plans (Fernando)
  3. Cathode redesign Bill's changes (Bill, Roger)
  4. Cosmic tests (Beni)
  5. Other
    • summer students, techs from FEL working on FDC project
    • talks at the Collaboration meeting FDC update
    • no meeting next week


Participants: Bill, Casey, Roger, Eugene, Glenn, Mark, Brian, Fernando, Simon, Beni, Kim, Lubomir


- We started our discussions from the last point: what work do we have for the summer students. There are several projects related to the UVA equipment for wire stringing: setting up the position measuring system (camera with computer), help in the development of a system for tension measurements. Not exactly related to that, but very important, Bill, Brian and others started discussion what else is needed to use the UVA equipment. We will have more discussions on the next meeting, but for now we will contact UVA to bring the equipment at JLab and store it in room 117.

- There's a possibility to have techs/engineers matrixed from FEL to work on the FDC project. Roger made the point that whenever FEL finds funding for them, they will abandon our project. We have to make sure that their contracts have the clause that they first have to finish their responsibilities for us.

- everybody is encouraged the look at the FDC Collaboration talks linked above

- because of the Collaboration meeting next week we will have no FDC meeting


- The plan is to have the space available starting Sep.1. Bill talked to Rusty and he has contacted two vendors for the clean room and will have a teleconference with them. We discussed also what can be done in 727 B.C. without the clean room if the space there is available before Sep.1.

- Brian and Mark made estimates for the manpower needed for the different steps in the FDC production. Bill will review the procedures and the numbers. The estimates for some of the procedures are not so certain, either they have never been done or the experience is not so solid.


- Fernando reduced the gain on the new cathode cards from 3.2mV/fC to 2.6mV/fC and the oscillations disappeared. The noise level of the cards when looking at the output of the shaper board (low amplification) is 10-15mV peak to peak. The oscillations are explained as a result of the increased capacitance of the system. We may have to reduce the amplification further for the 6-layer configuration. Bill asked if one can simulate/predict the amplification that works for the final configuration. Fernando: it will be difficult, effects are not linear.

- There were very interesting proposals how to improve the grounding to alleviate the oscillation problem. Fernando proposed to separate the grounds of the different layers; now one ground foil acts as a ground of two layers. We discussed how to increase to area/number of the connections between the ground and cathode foils and also to use the ground foil/copper tape to shield the input of the pre-amps. Eugene proposed to add thicker aluminum foil ring on the top of the ground plane covering only the frame area. Certainly there are good ideas; Bill has to look what modifications can be done to improve the grounding.

- Kim got a corrected estimate from CEM when using g10 (instead of FR4): $400 more. Bill stated, the flatness they can hold +/- 0.004" is acceptable for us. No quotes yet from TriLabs. Kim was having problems creating the step files for the PCBs that Bill needed for the 3D model. Fernando and Bill think it is important, so Fernando will help Kim, it may take a week.

- Gerard Visser will be here next week. Fernando plans for Thursday (05/13) to have him working on the FDC prototype to check the fADC125. The module will stay with us, unless Dave Abbott needs it, and it will be very important to start working with it.

Cathode redesign

Bill suggested moving the connector and opening for the rigid-flex outwards as explained in the above document. Roger will do this correction.

Cosmic tests

- On Tuesday the 3-layer prototype was installed back at the stand for testing with cosmics.

- Beni: additional delays are needed for the cathode signals to have them within the ADC gate. Beni will use long flat cables for that. He is working on his code to reconstruct the hit position from the cathode strips above and below the wire plane. The cathode pedestals are wide: RMS of ~80 ADC channels. Fernando suspects the post-amplifiers are noisy. He suggested using the shaper board instead of the post-amps and look for change of the widths.